OpenStudy (kidrah69):

Factor: 6x^2 + 11x - 10 I know how to do the problem but i cant think of factors.

OpenStudy (hysusonic):

what multiplys to -10 but adds up to 11?

OpenStudy (kidrah69):

-60 and 11

OpenStudy (hysusonic):

multiplies*

OpenStudy (kidrah69):

thats not right...

mathslover (mathslover):

Nope, you forgot that the coefficient of x^2 isn't 1. @hysusonic

OpenStudy (kidrah69):

we multiply 1 and c

OpenStudy (kidrah69):

*a

OpenStudy (hysusonic):

oh yea my bad @mathslover forgot about that

mathslover (mathslover):

Either you divide by 6, and make the coefficient of x^2 as 1. And then move on. But I usually prefer to follow quadratic formula and get the roots from there. And simply follow : \((x - \alpha) (x-\beta) \) .. .where \(\alpha\) and \(\beta\) are the roots.

OpenStudy (jhannybean):

\[6x^2+11x-10 \implies x^2+11x-60=0\]which two numbers multiply to give -60 and add 11?

OpenStudy (amistre64):

actually, what factors of -60 add up to 11 ?

OpenStudy (amistre64):

1 60 2 30 4 15 hmm, that looks good

OpenStudy (amistre64):

-4 and 15

OpenStudy (kidrah69):

ah ok that helps :) thanks i can do the rest of the problems myself just needed the factors.....thanks all! :)

OpenStudy (jhannybean):

Oh okay :)

mathslover (mathslover):

like here, \(6 \left(x^2 + \frac{11}{6} x - \frac{10}{6} \right) \) Find the roots of the equation in brackets. Roots are : 2/3 and -5/2 So, you can write it as : \(6 \left( x - \cfrac{2}{3}\right) \left(x + \cfrac{5}{2} \right) \) Which further gives : \((6x - 4)(6x + 15) \) So, roots are 4 and -15 !

OpenStudy (amistre64):

those are almost factors, they are useful in finding the rest of it

OpenStudy (amistre64):

mathslover explained it better than i would have lol

OpenStudy (amistre64):

(x - 4/6) (x + 15/6) simplify (x - 2/3) (x + 5/2) if theres still a bottom, put it back in front, thats where we got the 6 from (3x - 2) (2x+5)

mathslover (mathslover):

Yup!

mathslover (mathslover):

I wonder the roots are 4 and -15 and not -4 and 15. Or have I done something wrong?

OpenStudy (kidrah69):

15 and -4 are the roots... u have it right :P

OpenStudy (jhannybean):

\[\left(x-\frac{4}{6}\right)\left(x+\frac{15}{6}\right)=0\]reduce the fractions, and multiply diagonally whatever is unreducable. \[\left(x-\frac{2}{3}\right)\left(x+\frac{5}{2}\right)=0\]\[(3x-2)(2x+5)=0\]And then fromt here you just solve for x.

OpenStudy (jhannybean):

Oh, haha *highfive @amistre64 *

OpenStudy (kidrah69):

xD thanks guys! :D